Let f(x)=x^2+bx+c. It has turned out that equation f(x) = 2x - 7 has exactly one solution, and equation f(x)=21-6x also has exactly one solution. Find the largest value of parameter p such that equation f(x) has exactly one solution. 


1
Expert's answer
2021-10-06T17:04:11-0400

Given question is missing some parts and parameter p. We assume it as follows:

Let "f(x)=px^2+bx+c" . It has turned out that equation g(b) = 2b - 7 has exactly one solution, and equation h(c)=21-6c also has exactly one solution. Find the largest value of parameter p such that equation f(x) has exactly one solution. 

Solution:

"g(b)=2b-7=0
\\ \Rightarrow 2b=7
\\ \Rightarrow b=3.5"

"h(c)=21-6c=0
 \\\Rightarrow 6c=21
\\ \Rightarrow c=21/6=3.5"

So, "f(x)=px^2+3.5x+3.5"

Now f(x) has exactly one solution, so its discriminant D must be 0.

"D=B^2-4AC=0
\\ \Rightarrow 3.5^2-4p(3.5)=0
\\ \Rightarrow 12.25=14p
\\ \Rightarrow p=\dfrac{12.25}{14}
\\ \Rightarrow p=0.875"